OpenStudy (anonymous):

What is the value of x in the equation 2(2x − 5 − 4) = 160 − 17? 40.25 62.25 80.25 143.0

OpenStudy (anonymous):

@DIDITHEDAD @joojoe @MathMusician

OpenStudy (anonymous):

Distribute AGAIN lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

gimme a sec could you write the like formula? lol

OpenStudy (anonymous):

And subtract the side on the right of the equal sign

OpenStudy (anonymous):

its x2 right?

OpenStudy (anonymous):

2 * 2x 2 * -4 2 * -5

OpenStudy (anonymous):

ok yea i was right lol

OpenStudy (anonymous):

so 4x -8 -10

OpenStudy (anonymous):

okay.... 4x - 8 - 10 = 160 - 17 Try to do it by yourself, I have to go now... what is -8 + -10? -18 4x - 18 = 143 Now just add 18 and then divide by 4

OpenStudy (anonymous):

lol ok so then 143 + 18?

OpenStudy (anonymous):

@joojoe

OpenStudy (anonymous):

ohh so 161 divided by 4 = 40.25

OpenStudy (anonymous):

so a

OpenStudy (anonymous):

You are correct....sorry I just logged on real fast before I leave the house...nice job!

OpenStudy (anonymous):

lolthx